OpenStudy (anonymous):

sqrt3(sqrt6+sqrt3)+(sqrt8-5) sqrt3(sqrt6+sqrt3)+(sqrt8-5) @Mathematics

OpenStudy (anonymous):

3sqrt(2) + 3 + sqrt(3)

jhonyy9 (jhonyy9):

sqrt18 +sqrt9 +sqrt3 3sqrt2 +3 + sqrt3

OpenStudy (anonymous):

\[\sqrt{3}(\sqrt{6}+\sqrt{3})+(\sqrt{8}-5)\]

OpenStudy (anonymous):

neither of those are in my answers.

jhonyy9 (jhonyy9):

so than 3sqrt2 +3 +2-5 3sqrt2 +3-3 3sqrt2

OpenStudy (anonymous):

thats not one of my answers either.

jhonyy9 (jhonyy9):

why ?

OpenStudy (anonymous):

i got it nvm.

jhonyy9 (jhonyy9):

sqrt18=3sqrt2

OpenStudy (anonymous):

its 5sqrt2-2

myininaya (myininaya):

\[\sqrt{3}(\sqrt{6}+\sqrt{3})+(\sqrt{8}-5)=\sqrt{18}+\sqrt{9}+\sqrt{8}-5=3\sqrt{2}+3+2\sqrt{2}-5\] \[5 \sqrt{2}-2\]

OpenStudy (anonymous):

3sqrt(2) + 3 + 2sqrt(2) -5 5sqrt(2) -2